REMPI AND ZEKE STUDIES OF METHYLENE RADICAL.

\underline{SHILIANG WANG}, ZYGMUNT J. JAKUBEK, BENOIT SIMARD and PHIL R. BUNKER, Steacie Institute for Molecular Sciences, National Research Council, Ottawa, ON, Canada K1A 0R6.

The 3p and 4d Rydberg states of CH_{2} have been reinvestigated by resonance enhanced multiphoton ionization (REMPI) spectroscopy. Methylene radical has been produced by photolysis of ketene. In addition, preliminary results of zerokinetic-energy (ZEKE) photoelectron spectroscopy study of the ground state of CH_{2}^{+} will be presented.